OpenStudy (anonymous):

A sound wave in air has a speed of 340 m/s. If this sound wave is known to have a wavelength of 20 meters, what is its frequency?

OpenStudy (ash2326):

wavelength=20 meters speed= 340m/s frequency=speed/wavelength=340/20=17 Hz

OpenStudy (anonymous):

thanks what formula dis you use?

OpenStudy (ash2326):

frequency= speed/ wavelength

OpenStudy (anonymous):

Arrange the following sound waves according to speed from the slowest to the fastest: The sound waves in steel, the sound waves in cold air, the sound waves in water, and the sound waves in hot air.

OpenStudy (anonymous):

\[\lambda = {v \over f}\]This is the formula ash is using.

OpenStudy (anonymous):

Regarding your second question. Sound waves travel faster through denser mediums.